OpenStudy (anonymous):

if x varies inversely as v, and x = 35 when v = 3, find x when v=21

OpenStudy (anonymous):

x= 35 when you say " x inversely as v" it looks like this x = k/v wherein k is constant

OpenStudy (anonymous):

since x = 35 and v = 3 we can use the x = k/v

OpenStudy (anonymous):

\[35 = k/3\]

OpenStudy (anonymous):

so your k =?

OpenStudy (anonymous):

@catherinek

OpenStudy (anonymous):

k = 105

OpenStudy (anonymous):

so when v = 21 k =vx 105 = 21(x)

OpenStudy (anonymous):

so your x would be equal to?

OpenStudy (anonymous):

@catherinek are you still there? :)